OpenStudy (anonymous):

What is the maximum (highest) point on the graaph of this quadratic function: f(x)=-2x^2+8x+10 ? Help Please

OpenStudy (tkhunny):

Even seen "-b/(2a)"?

OpenStudy (anonymous):

Yea. Isn't that how to get the vertex? I think?

OpenStudy (tkhunny):

Perfect. That IS how to find the vertex. The vertex is the highest or lowest point that you seek.

OpenStudy (anonymous):

y = -2x^2 + 8x + 10 you could also try the derivative to find the point y'= -4x + 8 then set y' to 0 4x = 8 x=2 then substitute to the quadratic y = -2(2^2) + 8(2) + 10 y = 18 since y" is negative we could conclude that (2, 18) is a maxima kindly correct me if i missed or i was wrong at something

OpenStudy (anonymous):

It's okay thanks. I'm barely starting to learn to do this but yea I get it in a way.

OpenStudy (tkhunny):

Notice also how from y = -2x^2 + 8x + 10 we have -8/(2(-2)) = -8/(-4) = 2 and we have managed to use the calculus to find the same result.